You decide to buy 900 shares of stock at a price of $78 and an initial margin of 70 percent. What is the maximum percentage decline in the stock price before you will receive a margin call if the maintenance margin is 40 percent? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) price decline
Answer #1

Total investment in the shares =900*$78

=$70,200

100%-70% = 30% will be on margin( 70% of $70,200 is payed)

30% of $70,200( borrowed amount)= $21,060

stock price when before you receive a margin call

21,060/(1-maintenance margin)= 21,060/(1-0.40)

= $35,100

stock price =35100/900= $39

% decline in the stock price= 78-39/78 = 50%